# Tractable Metric Spaces and the Continuity of Magnitude

### v1

@@ Abstract
Magnitude is an isometric invariant of metric spaces introduced by Leinster. Since its inception, it has inspired active research into its connections with integral geometry, geometric measure theory, fractal dimensions, persistent homology, and applications in machine learning. In particular, when it comes to applications, continuity and stability of invariants play an important role. Although it has been shown that magnitude is nowhere continuous on the Gromov–Hausdorff space of finite metric spaces, positive results are possible if we restrict the ambient space. In this paper, we introduce the notion of tractable metric spaces, provide a characterization of these spaces, and establish several continuity results for magnitude in this setting. As a consequence, we offer a new proof of a known result stating that magnitude is continuous on the space of compact subsets of $\mathbb{R}$ with respect to the Hausdorff metric. Furthermore, we show that the magnitude function is Lipschitz when restricted to bounded subspaces of $\mathbb{R}$.
@@ 1 Introduction
Magnitude has been shown to encode many invariants from integral geometry and geometric measure theory, including volume, capacity, dimension, and intrinsic volumes [7, 11, 19, 6]. It has also been applied in other contexts—for example, to quantify biodiversity [17, 4], in machine learning, to measure the intrinsic diversity of latent representations [9] and in connection with persistent homology [13, 12, 2]. While the question of continuity of magnitude is mathematically interesting in its own right [5, 10], it becomes especially crucial in the context of data analysis [3], where a fundamental principle is that a good invariant should depend continuously on the input data.
@@
To this end, we introduce the class of tractable metric spaces. These are positive definite, Heine–Borel metric spaces in which all closed balls have finite magnitude. Examples include $\ell_{p}^{N}$ for $1\leq p\leq 2$, finite-dimensional vector subspaces of $L^{1}$, and their closed subspaces. We demonstrate that tractable metric spaces have properties that are naturally connected with the question of continuity of magnitude. For example, in a locally compact metric space, if the magnitude is finite and continuous at some compact subspace, that subspace is necessarily contained in some tractable neighborhood. Among others, we also establish the following equivalence: the continuity of magnitude on the space of compact subspaces of a tractable metric space $M$ is equivalent to the uniform continuity of magnitude on the space of finite subspaces of every bounded subspace $B\subseteq M$, and, in turn, to uniform continuity on the space of finite subspaces within each closed ball $\overline{B}_{M}\!\left(x,r\right)\subseteq M$. We also explore the one-point property [15, 8] in the context of tractable spaces. In particular, in a tractable star-shaped subset of a normed vector space, the magnitude function has the one-point property if and only if it is continuous at the star center. We also show that in any tractable normed vector space, magnitude is continuous at every singleton.
These results build a framework for deriving continuity properties of magnitude in concrete settings. For example, we can deduce that magnitude is continuous on the class of compact metric subspaces of $\mathbb{R}$ equipped with the Hausdorff metric (this statement has already appeared in the literature [5]). We also prove that magnitude is Lipschitz continuous when restricted to compact subsets of bounded subspaces of $\mathbb{R}$ (Theorem 4.6).
@@ 1.1 Notation
We denote the set of natural numbers by $\mathbb{N}$. We treat 00 as a natural number, so =$\mathbb{N}=\left\{{0,1,2,3,\ldots}\right\}$.
We denote the set of real numbers by $\mathbb{R}$, and the set of extended real numbers (with infinities included) by $\overline{\mathbb{R}}$. That is, $\overline{\mathbb{R}}=\mathbb{R}\cup\left\{{-\infty,\infty}\right\}$.
@@ 2.1 The Magnitude of a Matrix
A matrix may or may not have a weighting, and independently may or may not have a coweighting (for example, $\begin{pmatrix}1&0\\ 1&0\end{pmatrix}$ has infinitely many weightings, but no coweightings). However, a standard argument shows that if a matrix $A$ has at least one weighting and at least one coweighting, then the sum of components is the same for all weightings and all coweightings [7]. This sum is called the magnitude of $A$.
@@
  • The case $n=0$ is allowed; the unique $(0\times 0)$-matrix (the “empty matrix”) is positive definite, its determinant is $1$, its adjugate and inverse are again the empty matrix, and its magnitude is a sum of zero summands, so 00.
@@ 2.2 The Magnitude of Finite Metric Spaces
The magnitude of the empty space is 00, the magnitude of a singleton is $1$, and for two points at distance $u\in\mathbb{R}_{>0}$, the magnitude is $\frac{2}{1+e^{-u}}=1+\tanh(\frac{u}{2})$ where $\tanh$ denotes the hyperbolic tangent [7]. More generally, if $a_{1}<a_{2}<\ldots<a_{n}$ are real numbers (equipped with the usual metric), then $\mathsf{mag}(\left\{{a_{1},\ldots,a_{n}}\right\})=1+\sum_{i\in{\mathbb{N}}_{[1,n)}}\tanh(\frac{a_{i+1}-a_{i}}{2})$ [5].
However, we are rarely interested in magnitude of a single metric space at a time. More commonly studied (and much more informative) is the so-called magnitude function. Given a metric space $M$ and $t\in\mathbb{R}_{>0}$, the metric space $t\;\!M$ is defined to have the same underlying set as $M$, but all distances are multiplied by $t$. The magnitude function of a finite metric space $M$ is defined to be the partial map $\mathbb{R}_{>0}\mathrel{\rightharpoonup}\mathbb{R}$, $t\mapsto\mathsf{mag}(t\;\!M)$ [7]. It turns out that for every finite metric space, its magnitude function is defined for all but finitely many values, and is continuous on its domain of definition.
The definition of magnitude readily generalizes to pseudometric space (where two different points can be at zero distance). In this sense, the magnitude function can be extended to $t=0$. It is easy to calculate that for any non-empty finite metric space $M$ we have $\mathsf{mag}(0\;\!M)=1$. This extended magnitude function is no longer necessarily continuous, however; in fact, the discontinuity can be arbitrarily bad [15]. A non-empty finite metric space $M$ is said to have the one-point property (since $0\;\!M$ looks like one point) when its magnitude function is continuous at $t=0$, i.e. when $\lim_{t\to 0}\mathsf{mag}(t\;\!M)=1$. Which spaces have the one-point property is still a topic of ongoing research.
@@
Magnitude for finite subspaces can be seen as a generalization of the magnitude function. Let $M$ be a finite metric space with metric $d$. If $M=\emptyset$, then its magnitude function is constantly 00. Otherwise, fix $a\in M$ and define the metric on $M\times\mathbb{R}_{>0}$ as follows:
$\big{(}(x,t),(y,u)\big{)}\mapsto\begin{cases}t\cdot d(x,y)&\text{if $t=u$},\\ t\cdot d(x,a)+u\cdot d(y,a)+|t-u|&\text{if $t\neq u$}.\end{cases}$
Then the magnitude function of $M$ is a restriction of $\mathsf{mag}_{\mathrm{Fin}_{+}(M\times\mathbb{R}_{>0})}$, in the sense that $\mathsf{mag}(t\;\!M)=\mathsf{mag}_{\mathrm{Fin}_{+}(M\times\mathbb{R}_{>0})}(M\times\left\{{t}\right\})$. If $M$ is a finite subspace of a normed vector space $V$, we do not even need a product space, we simply have $\mathsf{mag}(t\;\!M)=\mathsf{mag}_{\mathrm{Fin}_{+}(V)}(t\cdot M)$.
@@ 2.3 Positive Definite Metric Spaces
A standard example of a positive definite space is $L_{1}$ [10]. Consequently, every metric space which isometrically embeds into $L_{1}$ is positive definite as well. This includes finite-dimensional $\ell_{p}^{N}$-spaces for $p\in{\mathbb{R}}_{[1,2]}$ [10].
Magnitude is well behaved for positive definite metric spaces [10]. To begin with, if $M$ is positive definite, then its magnitude $\mathsf{mag}_{\mathrm{Fin}(M)}\colon\mathrm{Fin}(M)\to\mathbb{R}$ is a total map. Additionally, it satisfies the following important property: for all $F^{\prime},F^{\prime\prime}\in\mathrm{Fin}(M)$, if $F^{\prime}\subseteq F^{\prime\prime}$, then $\mathsf{mag}_{\mathrm{Fin}(M)}(F^{\prime})\leq\mathsf{mag}_{\mathrm{Fin}(M)}(F^{\prime\prime})$. We say that positive metric spaces have inclusion-monotone magnitude. For the proof of this fact, see [7].
@@ 2.4 The Magnitude of Compact Metric Spaces
In principle, there is nothing stopping us to take this as the definition of the magnitude for any metric compact, but for general (non-positive definite) spaces, this is poorly behaved. In particular, for a finite metric space which is not positive definite, Definitions 2.3 and 2.7 might not yield the same magnitude [10].
@@
Let $a,b\in\mathbb{R}$, $a\leq b$. The magnitude of the interval ${\mathbb{R}}_{[a,b]}$ [6] with respect to the standard Euclidean metric is given by
@@
More generally [6], if $A=\bigcup_{i\in{\mathbb{N}}_{[1,n]}}{\mathbb{R}}_{[a_{i},b_{i}]}\subseteq\mathbb{R}$ is a finite union of closed bounded intervals such that $a_{i}\leq b_{i}<a_{i+1}\leq b_{i+1}$ for all $i\in{\mathbb{N}}_{[1,n)}$, then the magnitude is given by
@@
However, a more general question remains open as to whether restrictions of $\mathsf{mag}_{\mathrm{Cmp}(L_{1})}$ to finite-dimensional subspaces are continuous [3, Conjecture 1.3]. One of the purposes of this paper is to contribute to the answer.
@@ 3 Tractable Spaces
First, let us recall a weaker notion. In every metric space, every compact subspace is closed and bounded. Recall that a metric space is called Heine-Borel when the converse also holds, i.e. when its every closed bounded subspace is compact. This is equivalent to every closed ball being compact [14].
@@
    Take any $K\in\mathrm{Cmp}_{+}(M)$ and $r\in\mathbb{R}_{>0}$. Then for every $L\in\overline{B}_{\mathrm{Cmp}_{+}(M)}\!\left(K,r\right)$ we have $L\subseteq\overline{\mathrm{th}}_{M}(K,r)$. Since $\overline{\mathrm{th}}_{M}(K,r)$ is compact by the previous item, the space $\mathrm{Cmp}_{+}\big{(}\overline{\mathrm{th}}_{M}(K,r)\big{)}$ is compact by [1]. Hence, so is its closed subspace $\overline{B}_{\mathrm{Cmp}_{+}(M)}\!\left(K,r\right)$.
@@
Throughout the proof, we use the fact that a real normed vector space is Heine-Borel if and only if it is finite dimensional if and only if it is locally compact [16] (and if a metric space is locally compact, so is its isometric copy).
@@
    We can write $V=\bigcup_{n\in\mathbb{N}}\overline{B}_{V}\!\left(0,n\right)$. Every compact metric space is separable, so as a countable union of such spaces, $V$ is separable as well. Hence, by [10, Corollary 3.5], $V$ isometrically embeds into $L_{1}$.
@@
    Since $V$ is finite-dimensional, it is locally compact and separable. Since it isometrically embeds into $L_{1}$, it is positive definite. Hence, by [10, Corollary 3.5], there is a vector subspace $V^{\prime}\subseteq L_{1}$ which is an isometric copy of $V$. In particular, $V^{\prime}$ is locally compact, so finite dimensional.
@@
    (by inclusion-monotonicity of magnitude and the fact that $\delta\cdot\overline{B}_{S}\!\left(0,1\right)\subseteq\overline{B}_{S}\!\left(0,\delta\right)$ because $S$ is star-shaped at 00). Suppose $K\in\mathrm{Cmp}_{+}(S)$ (so there is some $a\in K$) and $d_{H}(\left\{{0}\right\},K)\leq\delta$. Then $\left\{{a}\right\}\subseteq K\subseteq\overline{B}_{S}\!\left(0,\delta\right)$, so $1\leq\mathsf{mag}_{\mathrm{Cmp}_{+}(S)}(K)\leq 1+\epsilon$. Hence $\big{|}\mathsf{mag}_{\mathrm{Cmp}_{+}(S)}(\left\{{0}\right\})-\mathsf{mag}_{\mathrm{Cmp}_{+}(S)}(K)\big{|}=\big{|}1-\mathsf{mag}_{\mathrm{Cmp}_{+}(S)}(K)\big{|}\leq\epsilon$.
@@
    Take any $K\in\mathrm{Cmp}_{+}(V)$ and $\epsilon\in\mathbb{R}_{>0}$. We have some $R\in\mathbb{R}_{>0}$ such that $K\subseteq\overline{B}_{S}\!\left(0,R\right)$. By assumption, there exists $\delta\in\mathbb{R}_{>0}$ such that for every $L\in\mathrm{Cmp}_{+}(S)$, if $d_{H}(\left\{{0}\right\},L)\leq\delta$, then $\big{|}\mathsf{mag}_{\mathrm{Cmp}_{+}(S)}(\left\{{0}\right\})-\mathsf{mag}_{\mathrm{Cmp}_{+}(S)}(L)\big{|}\leq\epsilon$. Hence, for $t\in{\mathbb{R}}_{(0,\min\left\{{\frac{\delta}{R},1}\right\}]}$, we have ($t\cdot K\subseteq S$ because $S$ is star-shaped at 00, and) $t\cdot K\subseteq\overline{B}_{S}\!\left(0,\delta\right)$, so $d_{H}(\left\{{0}\right\},t\cdot K)\leq\delta$, therefore $\mathsf{mag}_{\mathrm{Cmp}_{+}(S)}(t\cdot K)\leq\mathsf{mag}_{\mathrm{Cmp}_{+}(S)}(\left\{{0}\right\})+\epsilon=1+\epsilon$.

### v2

# arXiv:2506.21128v2[math.GN] 02 Mar 2026

@@ Abstract
Magnitude is an isometric invariant of metric spaces introduced by Leinster. Since its inception, it has inspired active research into its connections with integral geometry, geometric measure theory, fractal dimensions, persistent homology, and applications in machine learning. In particular, when it comes to applications, continuity and stability of invariants play an important role. Although it has been shown that magnitude is nowhere continuous on the Gromov–Hausdorff space of finite metric spaces, positive results are possible if we restrict the ambient space. In this paper, we introduce the notion of tractable metric spaces, provide a characterization of these spaces, and establish several continuity results for magnitude in this setting. As a consequence, we offer a new proof of a known result stating that magnitude is continuous on the space of compact subsets of $\mathbb{R}$ with respect to the Hausdorff metric. Furthermore, we show that magnitude is Lipschitz continuous when restricted to bounded subspaces of $\mathbb{R}$.
@@ 1 Introduction
Magnitude has been shown to encode many invariants from integral geometry and geometric measure theory, including volume, capacity, dimension, and intrinsic volumes [9, 11, 24, 7]. It has also been applied in other contexts — for example, to quantify biodiversity [22, 5], in machine learning, to measure the intrinsic diversity of latent representations [10] and in connection with persistent homology [16, 15, 3]. While the question of continuity of magnitude is mathematically interesting in its own right [6, 12, 20, 21], it becomes especially crucial in the context of data analysis [4], where a fundamental principle is that a good invariant should depend continuously on the input data.
@@
To this end, we introduce the class of tractable metric spaces. These are positive definite, Heine–Borel metric spaces in which all closed balls have finite magnitude. Examples include $\ell_{p}^{N}$ for $1\leq p\leq 2$, finite-dimensional vector subspaces of $L^{1}$, and their closed subspaces. We demonstrate that tractable metric spaces have properties that are naturally connected with the question of continuity of magnitude. For example, in a locally compact metric space, if the magnitude is finite and continuous at some compact subspace, that subspace is necessarily contained in some tractable neighborhood. Among others, we also establish the following equivalence: the continuity of magnitude on the space of compact subspaces of a tractable metric space $M$ is equivalent to the uniform continuity of magnitude on the space of finite subspaces of every bounded subspace $B\subseteq M$, and, in turn, to uniform continuity on the space of finite subspaces within each closed ball $\overline{B}_{M}\!\left(x,r\right)\subseteq M$. We also explore the one-point property [18, 8] in the context of tractable spaces. In particular, in a tractable star-shaped subset of a normed vector space, magnitude of compact/finite subsets is continuous at the star center if and only if the magnitude function of every non-empty compact/finite subset has the one-point property. We also show that in any tractable normed vector space, magnitude is continuous at every singleton.
These results build a framework for deriving continuity properties of magnitude in concrete settings. For example, we can deduce that magnitude is continuous on the class of compact metric subspaces of $\mathbb{R}$ equipped with the Hausdorff metric (this statement has already appeared in the literature, see [6, Corollary 5.3.13] and [12, Corollary 2.12]). We also prove that magnitude is Lipschitz continuous when restricted to compact subsets of bounded subspaces of $\mathbb{R}$ (Theorem 4.6).
@@ 1.1 Notation
We denote the set of natural numbers by $\mathbb{N}$. We treat 0 as a natural number, so $\mathbb{N}=\left\{{0,1,2,3,\ldots}\right\}$.
[…]
@@ 2.1 The Magnitude of a Matrix
A matrix may or may not have a weighting, and independently may or may not have a coweighting (for example, $\begin{pmatrix}1&0\\ 1&0\end{pmatrix}$ has infinitely many weightings, but no coweightings). However, a standard argument shows that if a matrix $A$ has at least one weighting and at least one coweighting, then the sum of components is the same for all weightings and all coweightings [9, Lemma 1.1.2]. This sum is called the magnitude of $A$.
@@
  • The case $n=0$ is allowed; the unique $(0\times 0)$-matrix (the “empty matrix”) is positive definite, its determinant is $1$, its adjugate and inverse are again the empty matrix, and its magnitude is a sum of zero summands, so 0.
@@ 2.2 The Magnitude of Finite Metric Spaces
The magnitude of the empty space is 0, the magnitude of a singleton is $1$, and for two points at distance $u\in\mathbb{R}_{>0}$, the magnitude is $\frac{2}{1+e^{-u}}=1+\tanh(\frac{u}{2})$ where $\tanh$ denotes the hyperbolic tangent [9, Examples 2.1.1]. More generally, if $a_{1}<a_{2}<\ldots<a_{n}$ are real numbers (equipped with the usual metric), then $\mathsf{mag}(\left\{{a_{1},\ldots,a_{n}}\right\})=1+\sum_{i\in{\mathbb{N}}_{[1,n)}}\tanh(\frac{a_{i+1}-a_{i}}{2})$ [9, Corollary 2.3.4].
However, we are rarely interested in magnitude of a single metric space at a time. More commonly studied (and much more informative) is the so-called magnitude function. Given a metric space $M$ and $t\in\mathbb{R}_{>0}$, the metric space tMt\;\!M is defined to have the same underlying set as $M$, but all distances are multiplied by $t$. The magnitude function of a finite metric space $M$ is defined to be the partial map $\mathbb{R}_{>0}\mathrel{\rightharpoonup}\mathbb{R}$, $t\mapsto\mathsf{mag}(t\;\!M)$ [9]. It turns out that for every finite metric space, its magnitude function is defined for all but finitely many values, and is continuous on its domain of definition.
The definition of magnitude readily generalizes to pseudometric space (where two different points can be at zero distance). In this sense, the magnitude function can be extended to $t=0$. It is easy to calculate that for any non-empty finite metric space $M$ we have $\mathsf{mag}(0\;\!M)=1$. This extended magnitude function is no longer necessarily continuous, however; in fact, the discontinuity can be arbitrarily bad [18, Theorem 3.8]. A non-empty finite metric space $M$ is said to have the one-point property (since 0M0\;\!M looks like one point) when its magnitude function is continuous at $t=0$, i.e. when $\lim_{t\to 0}\mathsf{mag}(t\;\!M)=1$. Which spaces have the one-point property is still a topic of ongoing research.
@@
Magnitude for finite subspaces can be seen as a generalization of the magnitude function (aside from the trivial case of the empty space whose magnitude function is constantly 0), as we now demonstrate.
Proposition 2.5.
Let $M$ be a finite metric space with metric $d$ and let $a\in M$. Then
$\rho\big((x,t),(y,u)\big):=\begin{cases}t\cdot d(x,y)&\text{if $t=u$},\\ t\cdot d(x,a)+u\cdot d(y,a)+|t-u|&\text{if $t\neq u$}\end{cases}$
is a metric on $M\times\mathbb{R}_{>0}$. Moreover, for this metric, for every $t\in\mathbb{R}_{>0}$ the map $t\;\!M\to M\times\mathbb{R}_{>0}$, $x\mapsto(x,t)$ is an isometric embedding.
Proof.
We clearly have $\rho\big((x,t),(x,t)\big)=0$. Conversely, assume $\rho\big((x,t),(y,u)\big)=0$. If we had $t\neq u$, then $\rho\big((x,t),(y,u)\big)=t\cdot d(x,a)+u\cdot d(y,a)+|t-u|\geq|t-u|>0$, a contradiction, so $t=u$. Hence $0=\rho\big((x,t),(y,u)\big)=t\cdot d(x,y)$, and since $t>0$, we conclude $d(x,y)=0$. Thus $x=y$, so $(x,t)=(y,u)$.
Obviously, $\rho$ is symmetric. It remains to verify the triangle inequality. Take any $(x,t),(y,u),(z,v)\in M\times\mathbb{R}_{>0}$.
Case $t=u=v$:
$\rho\big((x,t),(y,u)\big)+\rho\big((y,u),(z,v)\big)=t\cdot d(x,y)+u\cdot d(y,z)=$
$t\cdot\big(d(x,y)+d(y,z)\big)\geq t\cdot d(x,z)=\rho\big((x,t),(z,v)\big).$
Case $t\neq u=v$:
$\rho\big((x,t),(y,u)\big)+\rho\big((y,u),(z,v)\big)=t\cdot d(x,a)+u\cdot d(y,a)+|t-u|+u\cdot d(y,z)=$
$t\cdot d(x,a)+u\cdot\big(d(y,a)+d(y,z)\big)+|t-u|\geq$
$\geq t\cdot d(x,a)+u\cdot d(z,a)+|t-u|=\rho\big((x,t),(z,v)\big).$
Case $t=u\neq v$ is treated in the same way.
Case $t\neq u\neq v$, $t=v$:
$\rho\big((x,t),(y,u)\big)+\rho\big((y,u),(z,v)\big)=$
$t\cdot d(x,a)+u\cdot d(y,a)+|t-u|+u\cdot d(y,a)+v\cdot d(z,a)+|u-v|\geq$
$\geq t\cdot\big(d(x,a)+d(z,a)\big)\geq t\cdot d(x,z)=\rho\big((x,t),(z,v)\big).$
Case $t\neq u\neq v$, $t\neq v$:
$\rho\big((x,t),(y,u)\big)+\rho\big((y,u),(z,v)\big)=$
$t\cdot d(x,a)+u\cdot d(y,a)+|t-u|+u\cdot d(y,a)+v\cdot d(z,a)+|u-v|\geq$
$\geq t\cdot d(x,a)+v\cdot d(z,a)+|t-v|=\rho\big((x,t),(z,v)\big).$
The fact that the map $t\;\!M\to M\times\mathbb{R}_{>0}$, $x\mapsto(x,t)$ is an isometric embedding is immediate from the definition of $\rho$. ∎
It follows from this proposition that the magnitude function of $M$ is a restriction of $\mathsf{mag}_{\mathrm{Fin}_{+}(M\times\mathbb{R}_{>0})}$, in the sense that $\mathsf{mag}(t\;\!M)=\mathsf{mag}_{\mathrm{Fin}_{+}(M\times\mathbb{R}_{>0})}(M\times\left\{{t}\right\})$. If $M$ is a finite subspace of a normed vector space $V$, we do not even need a product space, we simply have $\mathsf{mag}(t\;\!M)=\mathsf{mag}_{\mathrm{Fin}_{+}(V)}(t\cdot M)$.
@@ 2.3 Positive Definite Metric Spaces
A standard example of a positive definite space is $L_{1}$ [12, Theorem 3.6]. Consequently, every metric space which isometrically embeds into $L_{1}$ is positive definite as well. This includes $L_{p}$ and $\ell_{p}^{N}$-spaces for $p\in{\mathbb{R}}_{[1,2]}$ [2, Theorem 2.1].
Magnitude is well behaved for positive definite metric spaces [12]. To begin with, if $M$ is positive definite, then its magnitude $\mathsf{mag}_{\mathrm{Fin}(M)}\colon\mathrm{Fin}(M)\to\mathbb{R}$ is a total map. Additionally, it satisfies the following important property: for all $F^{\prime},F^{\prime\prime}\in\mathrm{Fin}(M)$, if $F^{\prime}\subseteq F^{\prime\prime}$, then $\mathsf{mag}_{\mathrm{Fin}(M)}(F^{\prime})\leq\mathsf{mag}_{\mathrm{Fin}(M)}(F^{\prime\prime})$. We say that positive metric spaces have inclusion-monotone magnitude. For the proof of this fact, see [9, Corollary 2.4.4].
@@ 2.4 The Magnitude of Compact Metric Spaces
In principle, there is nothing stopping us to take this as the definition of the magnitude for any metric compact, but for general (non-positive definite) spaces, this is poorly behaved. In particular, for a finite metric space which is not positive definite, Definitions 2.3 and 2.8 might not yield the same magnitude [12].
@@
Let $\in\mathbb{R}, aba\leq b$. The magnitude of the interval ${\mathbb{R}}_{[a,b]}$ [7, Theorem 7] with respect to the standard Euclidean metric is given by
@@
More generally [6, Corollary 5.4.3], if $A=\bigcup_{i\in{\mathbb{N}}_{[1,n]}}{\mathbb{R}}_{[a_{i},b_{i}]}\subseteq\mathbb{R}$ is a finite union of closed bounded intervals such that $a_{i}\leq b_{i}<a_{i+1}\leq b_{i+1}$ for all $i\in{\mathbb{N}}_{[1,n)}$, then the magnitude is given by
@@
However, a more general question remains open as to whether restrictions of $\mathsf{mag}_{\mathrm{Cmp}(L_{1})}$ to finite-dimensional subspaces are continuous [4, Conjecture 1.3]. One of the purposes of this paper is to contribute to the answer.
@@ 3 Tractable Spaces
First, let us recall a weaker notion. In every metric space, every compact subspace is closed and bounded. Recall that a metric space is called Heine-Borel (or proper) when the converse also holds, i.e. when its every closed bounded subspace is compact. This is equivalent to every closed ball being compact [17, 1 Definitions, Chapter H].
@@
    Take any $K\in\mathrm{Cmp}_{+}(M)$ and $r\in\mathbb{R}_{>0}$. Then for every $L\in\overline{B}_{\mathrm{Cmp}_{+}(M)}\!\left(K,r\right)$ we have $L\subseteq\overline{\mathrm{th}}_{M}(K,r)$. Since $\overline{\mathrm{th}}_{M}(K,r)$ is compact by the previous item, the space $\mathrm{Cmp}_{+}\big(\overline{\mathrm{th}}_{M}(K,r)\big)$ is compact by [1, Theorem 7.3.8]. Hence, so is its closed subspace $\overline{B}_{\mathrm{Cmp}_{+}(M)}\!\left(K,r\right)$.
@@
Throughout the proof, we use the fact that a real normed vector space is Heine-Borel if and only if it is finite dimensional if and only if it is locally compact [19, Theorems 1.22, 1.23] (and if a metric space is locally compact, so is its isometric copy).
@@
    We can write $V=\bigcup_{n\in\mathbb{N}}\overline{B}_{V}\!\left(0,n\right)$. Every compact metric space is separable, so as a countable union of such spaces, $V$ is separable as well. Hence, by [12, Corollary 3.5], $V$ isometrically embeds into $L_{1}$.
@@
    Since $V$ is finite-dimensional, it is locally compact and separable. Since it isometrically embeds into $L_{1}$, it is positive definite. Hence, by [12, Corollary 3.5], there is a vector subspace $V^{\prime}\subseteq L_{1}$ which is an isometric copy of $V$. In particular, $V^{\prime}$ is locally compact, so finite dimensional.
@@
As stated, the spaces $\ell_{p}^{N}$ are tractable for $p\in{\mathbb{R}}_{[1,2]}$. Recall that for $p\in{\mathbb{R}}_{(0,1)}$, the usual definition of pp-“norm” is not in fact a norm, but its $p$-th power does induce a metric on $\mathbb{R}^{N}$, denoted by $d_{p}$. These spaces also often appear in the context of magnitude, so we characterize their tractability here.
Proposition 3.6.
The following statements are equivalent for every $N\in\mathbb{N}$ and $p\in{\mathbb{R}}_{(0,1)}$.
  1. 1.
    The space $(\mathbb{R}^{N},d_{p})$ is tractable.
  2. 2.
    The space $(\mathbb{R}^{N},d_{p})$ is positive definite.
  3. 3.
    There exists a linear map $T\colon\mathbb{R}^{N}\to L_{p}({\mathbb{R}}_{[0,1]})$ such that $\|Tx\|_{p}=\|x\|$ for every $x\in\mathbb{R}^{N}$.
Proof.
The equivalence between (ii) and (iii) is a known result [6, Theorem 4.10]. Statement (i) implies (ii) by definition. It remains to verify that (ii) implies (i).
Note that every closed ball in $(\mathbb{R}^{N},d_{p})$ is also bounded and closed in, say, the $\infty$-metric, and the metrics $d_{p}$ and $d_{\infty}$ are equivalent. Hence all closed $p$-balls are compact. That they have finite magnitude follows from [6, Proposition 4.13]. ∎
@@
    (by inclusion-monotonicity of magnitude and the fact that $\delta\cdot\overline{B}_{S}\!\left(0,1\right)\subseteq\overline{B}_{S}\!\left(0,\delta\right)$ because $S$ is star-shaped at 0). Suppose $K\in\mathrm{Cmp}_{+}(S)$ (so there is some $a\in K$) and $d_{H}(\left\{{0}\right\},K)\leq\delta$. Then $\left\{{a}\right\}\subseteq K\subseteq\overline{B}_{S}\!\left(0,\delta\right)$, so $1\leq\mathsf{mag}_{\mathrm{Cmp}_{+}(S)}(K)\leq 1+\epsilon$. Hence $\big|\mathsf{mag}_{\mathrm{Cmp}_{+}(S)}(\left\{{0}\right\})-\mathsf{mag}_{\mathrm{Cmp}_{+}(S)}(K)\big|=\big|1-\mathsf{mag}_{\mathrm{Cmp}_{+}(S)}(K)\big|\leq\epsilon$.
@@
    Take any $K\in\mathrm{Cmp}_{+}(V)$ and $\epsilon\in\mathbb{R}_{>0}$. We have some $R\in\mathbb{R}_{>0}$ such that $K\subseteq\overline{B}_{S}\!\left(0,R\right)$. By assumption, there exists $\delta\in\mathbb{R}_{>0}$ such that for every $L\in\mathrm{Cmp}_{+}(S)$, if $d_{H}(\left\{{0}\right\},L)\leq\delta$, then $\big|\mathsf{mag}_{\mathrm{Cmp}_{+}(S)}(\left\{{0}\right\})-\mathsf{mag}_{\mathrm{Cmp}_{+}(S)}(L)\big|\leq\epsilon$. Hence, for $t\in{\mathbb{R}}_{(0,\min\left\{{\frac{\delta}{R},1}\right\}]}$, we have ($t\cdot K\subseteq S$ because $S$ is star-shaped at 0, and) $t\cdot K\subseteq\overline{B}_{S}\!\left(0,\delta\right)$, so $d_{H}(\left\{{0}\right\},t\cdot K)\leq\delta$, therefore $\mathsf{mag}_{\mathrm{Cmp}_{+}(S)}(t\cdot K)\leq\mathsf{mag}_{\mathrm{Cmp}_{+}(S)}(\left\{{0}\right\})+\epsilon=1+\epsilon$.